Racin' Scooters is introducing a new product and has an expected change in EBIT of $425,000. Racin' Scooters has a 33 percent marginal tax rate. The project will produce $90,000 of depreciation per year. In​ addition, the project will cause the following changes in year​ 1:

What is the​project's free cash flow in year​ 1?

   WITHOUT THE PROJECT   WITH THE PROJECT
Accounts receivable   44,000   62,000
Inventory   65,000   89,000
Accounts payable   75,000   93,000
Answer #1

Answer :project free cashflow in year 1 is $350,750

Explanation

Without the project with the project
Accounts receivable $44000 62000
Inventory $65000 89000
Accounts payable 75000 93000
Working capital =current assets -current liabilities 34000 58000

Net change in working capital =34000-58000=-24000

Project Free Cashflow =EBIT(1-T)+depreciation +change in net working capital -capital expenditure

=$425000(1-0.33)+90000-24000

=$28,4750+90000-24000

=$350,750(answer)